OpenStudy (anonymous):

at noon ship A is 40 miles due west of ship B. ship A is sailing west at 18mph and ship B is sailing north @ 23 mph. how fast is distance b/w the ships changing at 7pm

OpenStudy (anonymous):

distance = 40 + 7(18+23) = 40 +287 = 327 miles

OpenStudy (anonymous):

distance changing by 41miles/hour